OpenStudy (anonymous):

X to the power of 2=196 please explain

OpenStudy (tkhunny):

Explain what? Solve like any other quadratic equation.

OpenStudy (anonymous):

X to the power of 2=196

OpenStudy (anonymous):

It says find the equation for x

OpenStudy (tkhunny):

That makes no sense. You are given the equation. Perhaps find the solutions?

OpenStudy (anonymous):

Idk this is tooooo confusing

OpenStudy (tkhunny):

It's just a quadratic equation. Have you solved any of those?

OpenStudy (anonymous):

?????? Please help I put a pic of it

OpenStudy (anonymous):

And no

OpenStudy (tkhunny):

Yup. Solve the equation for x. You are given the equation. Your task is to find a value for x that satisfies the equation. 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 5^2 = 25 6^2 = 36 ... 10^2 = 100 ... 15^2 = 225 Are we getting close?

OpenStudy (anonymous):

Umm ain't 225 higher then 196

OpenStudy (tkhunny):

Exactly. So,I must have missed it. Can you find the right value?

OpenStudy (anonymous):

Ohh so it would be 14 to the power of 2 right

OpenStudy (tkhunny):

Done. Move on to the next one.

OpenStudy (anonymous):

Oh ok that all I have to do

OpenStudy (unklerhaukus):

what about the other solution?

OpenStudy (unklerhaukus):

\[x^2=196\\ x\times x = 196\\ x = \pm\sqrt{196}\]

OpenStudy (tkhunny):

It seemed to me that we were not worried about that other solution, based on the visual observation of the problem statement. We certainly aren't going to expect a student at this level to find all three solutions to the next problem.